Claude Crépeau

Curriculum Vitæ

Informations personnelles

citoyenneté: Canadien langues parlées, lues et écrites: français et anglais date de naissance: 17/2/62 lieu de naissance: Montréal, Québec

Adresses

Rés.: 280 rue Provencher, Mont St-Hilaire, Québec J3H 3M2 Canada Tél.: +1 (514) 446-9441 Bur.: Département d’Informatique et R.O., Université de Montréal, Tél.: +1 (514) 343-2446 C.P. 6128 succ. C-V, Montréal, Québec H3S 3J7 Canada FAX.: +1 (514) 343-5834 adresse électronique: crepeau@iro.umontreal.ca

Expérience de travail

institution  période

poste

Université de Montréal, Dép. IRO, Montréal, Québec  6/95–

Professeur Agrégé

Laboratoire d’Informatique de l’École Normale Supérieure, Paris, France  10/91–6/95
Chargé de recherches niveau 1 au C.N.R.S.  (titularisé 4/93)
Université de Montréal, Dép. IRO, Montréal, Québec  6/94–6/95

Professeur Associé

Université Paris–Sud, Laboratoire de Recherche en Informatique, Orsay, France  12/89–10/91

Maître de conférence associé/Chercheur postdoctoral

Massachusetts Institute of Technology, Cambridge MA, U.S.A.  9/86–12/89

Assistant de recherche

Université de Montréal, Dép. IRO, Montréal, Québec  9/85–5/86
Assistant d’enseignement  9/83 - 12/84
Bell-Nothern Research, Îles des Soeurs (Verdun), Québec  5/84–9/86

Membre du personel scientifique

Études

diplôme  période

institution

Ph.D. Computer Science (cryptographie)  9/86–12/89
dirigé par Silvio Micali  M. I. T.
Maîtrise ès Sciences Informatique (cryptographie)  9/84–8/86
dirigé par Gilles Brassard  Université de Montréal
participation au “Crypto course”  10/85
organisé par David Chaum  C.W.I.
Baccalauréat ès Sciences  9/81–5/84
Mathématiques–Informatique  Université de Montréal

Contributions à la communauté scientifique

contributions  période
membre du comité de programme  A95

du colloque “ACM Conference on Computer and Communications Security’

Associate Editor for Complexity and Cryptography  1/95–1/98

de l’“IEEE Transactions on Information Theory”

membre du comité éditorial du “Journal of Cryptology”  1/91–1/95
membre du comité de programme des colloques “Eurocrypt ’94” et  H94

“XIV International Conference of the Chilean Computer Science Society”

membre du comité de programme du colloque “Eurocrypt ’90”  H90
membre du comité de programme du colloque “Crypto ’89”  H89

Visites Scientifiques

institution  période
Université de Montréal, Dép. IRO, Montréal, Québec 1–12/94 6–9/93 6–12/92
NEC Research, South Brunswick NJ, U.S.A  3/92
CWI, Amsterdam, les Pays-Bas  12/91 6–7/87
Université d’Århus, Århus, Danemark  10–11/91 6/88
IBM, Almaden Research Center, San José CA, U.S.A.  8/88
Bell Communication Research, Morristown NJ, U.S.A.  8/87

Enseignement

institution  cours  niveau  durée  période
Asian Institute of Tech.  Cryptographie  30h  12/94
Université de Montréal  Info. Théorique (IFT2102)  B.Sc.  60h  A94
Cryptologie (IFT6180)  M.Sc.–Ph.D. 60h  A92 H94
École Normale Supérieure Protocoles Cryptographiques  D.E.A.  12h H92 H93 A93
École Nationale Supérieure Complexité et Cryptographie 3è année  21h H92 H93 A93
de Technologie Avancée  ingénieure
Université Paris–Sud  Calculabilité et Complexité  maîtrise  25h  H91 H92
Cryptographie  D.E.A.  30h  H90 H91
Algorithmique  M.I.A.G.  30h  H90

Direction d’étudiants

institution  étudiants  période
Université de Montréal, Dép. IRO, Montréal, Québec 2 M.Sc (co-dir.) 9/93–9/95
1 Ph.D. (co-dir.) 1/92–9/95
1 Ph.D. (co-dir.) 1/93–1/96
Université Paris-Sud, Orsay, France 1 D.E.A. 1/92–9/92
1 Doctorat  9/93–9/96

Publications

thèses et mémoires

[Cr90a] C. Crépeau. Correct and Private Reductions among Oblivious Transfers. PhD thesis, Department of Elec. Eng. and Computer Science, Massachusetts Institute of Technology, 1990. Supervised by Silvio Micali.

[Cr86a] C. Crépeau. Protocole cryptographique de poker à l’aveugle permettant la confidentialié de la stratégie. Master’s thesis, Département d’Informatique et de R.O., Université de Montréal, 1986. Supervised by Gilles Brassard.

revues scientifiques

[BBCM95] C.H. Bennett, G. Brassard, C. Crépeau, and U. Maurer. Generalized Privacy Amplification. IEEE Transaction on Information Theory, to appear, November 1995.

[Cr94] C. Crépeau. Quantum Oblivious Transfer. Journal of Modern Optics, special issue on Quantum Communication and Cryptography. Volume 41, Number 12, pp. 2445-2454, December 1994.

[BBCJPW93] C.H. Bennett, G. Brassard, C. Crépeau, R. Jozsa, A. Peres, and W. Wootters. Teleporting an unkown quantum state by dual classical and EPR channels. Physical Review Letter, 70:1895–1898, 1993.

[BCY90] G. Brassard, C. Crépeau, and M. Yung. Constant-round perfect zero-knowledge computationally convincing protocols. Theoretical Computer Science, 84:23–52, 1991.

[BBCGP8B] P. Beauchemin, G. Brassard, C. Crépeau, C. Goutier, and C. Pomerance. The generation of random numbers that are probably prime. Journal of Cryptology, 1(1):53–64, 1988.

[BCC88a] G. Brassard, D. Chaum, and C. Crépeau. An introduction to minimum disclosure. Quarterly, 1:3–17, 1988.

[BCC88b] G. Brassard, D. Chaum, and C. Crépeau. Minimum disclosure proofs of knowledge. JCSS, 37:156–189, 1988.

compte-rendus de colloques avec comité de lecture

[CGT95] C. Crépeau, J. van de Graaf and A. Tapp. Committed Oblivious Transfer and Private Multi-Party Computations. Advances in Cryptology: Proceedings of Crypto ’95, to appear, August 1995.

[CS95] C. Crépeau, and L. Salvail. Quantum Oblivious Mutual Identification. Advances in Cryptology: Proceedings of Eurocrypt ’95, Springer-Verlag, pages 133-147, 1995.

[BBCM94] C.H. Bennett, G. Brassard, C. Crépeau, and U. Maurer. Generalized Privacy Amplification (Abstract). Proceedings of the IEEE Symposium on Information Theory, page 350. IEEE, 1994.

[BCJL93] G. Brassard, C. Crépeau, R. Jozsa, and D. Langlois. A quantum bit commitment scheme provably unbreakable by both parties. In 29^{th} Symp. on Found. of Computer Sci., pages 42–52. IEEE, 1993.

[CK93] C. Crépeau and J. Kilian. Discreet solitary games. In Advances in Cryptology: Proceedings of Crypto ’93, Lecture Notes in Computer Science. Springer-Verlag, pages 319–330,1993.

[Cr92] C. Crépeau. Cryptographic primitives and quantum theory. In Proceedings of the second Physics of Computation Workshop, PhysComp ’92, pages 200–204, 1993.

[BBCS92] C.H. Bennett, G. Brassard, C. Crépeau, and M.-H. Skubiszewska. Practical quantum oblivious transfer protocols. In Advances in Cryptology: Proceedings of Crypto ’91, volume 576 of Lecture Notes in Computer Science, pages 351–366. Springer-Verlag, 1992.

[CS91b] C. Crépeau and M. Sántha. Efficient reductions among oblivious transfer protocols based on new self-intersecting codes. In Sequences II, Methods in Communications, Security, and Computer Science, pages 360–368. Springer-Verlag, 1991.

[BC91] G. Brassard and C. Crépeau. Quantum bit commitment and coin tossing protocols. In Advances in Cryptology: Proceedings of Crypto ’90, volume 537 of Lecture Notes in Computer Science, pages 49–61. Springer-Verlag, 1991.

[BCLL91] G. Brassard, C. Crépeau, S. Laplante, and C. Léger. Computationally convincing proofs of knowledge. In Proceedings of STACS ’91, volume 480 of Lecture Notes in Computer Science, pages 251–262. Springer-Verlag, 1991.

[CS91a] C. Crépeau and M. Sántha. On the reversibility of oblivious transfer. In Advances in Cryptology: Proceedings of Eurocrypt ’91, volume 547 of Lecture Notes in Computer Science, pages 106–113. Springer-Verlag, 1991.

[BC90] G. Brassard and C. Crépeau. Sorting out zero-knowledge. In Advances in Cryptology: Proceedings of Eurocrypt ’89, volume 434 of Lecture Notes in Computer Science, pages 150–154. Springer-Verlag, 1990.

[Cr90b] C. Crépeau. Verifiable disclosure of secrets and applications. In Advances in Cryptology: Proceedings of Eurocrypt ’89, volume 434 of Lecture Notes in Computer Science, pages 181–191. Springer-Verlag, 1990.

[BCY89] G. Brassard, C. Crépeau, and M. Yung. Everything in NP can be argued in perfect zeroknowledge in a constant number of rounds. In 16^{th} ICALP, volume 372 of Lecture Notes in Computer Science, pages 123–136. Springer-Verlag, 1989.

[CCD88] D. Chaum, C. Crépeau, and I. Damgård. Multiparty unconditionally secure protocols. In 19^{th} Symp. on Theory of Computing, pages 11–19. ACM, 1988.

[CK88] C. Crépeau and J. Kilian. Achieving oblivious transfer using weakened security assumptions. In 29^{th} Symp. on Found. of Computer Sci., pages 42–52. IEEE, 1988.

[CK90] C. Crépeau and J. Kilian. Weakening security assumptions and oblivious transfer. In S. Goldwasser, editor, Advances in Cryptology: Proceedings of Crypto ’88, volume 403 of Lecture Notes in Computer Science, pages 2–7. Springer-Verlag, 1990.

[Cr88] C. Crépeau. Equivalence between two flavours of oblivious transfers (abstract). In C. Pomerance, editor, Advances in Cryptology: Proceedings of Crypto ’87, volume 293 of Lecture Notes in Computer Science, pages 350–354. Springer-Verlag, 1988.

[BBCGP87] P. Beauchemin, G. Brassard, C. Crépeau, C. Goutier, and C. Pomerance. Two observations on probabilistic primality testing. In Advances in Cryptology: Proceedings of Crypto ’86, volume 263 of Lecture Notes in Computer Science, pages 443–450. Springer-Verlag, 1987.

[BC87] G. Brassard and C. Crépeau. Zero-knowledge simulation of boolean circuits (extended abstract). In Advances in Cryptology: Proceedings of Crypto ’86, volume 263 of Lecture Notes in Computer Science, pages 223–233. Springer-Verlag, 1987.

[BCR87] G. Brassard, C. Crépeau, and J.–M. Robert. All-or-nothing disclosure of secrets (extended abstract). In Advances in Cryptology: Proceedings of Crypto ’86, volume 263 of Lecture Notes in Computer Science, pages 234–238. Springer-Verlag, 1987.

[Cr87] C. Crépeau. A zero-knowledge poker protocol that achieves confidentiality of the players’ strategy or how to achieve an electronic poker face. In Advances in Cryptology: Proceedings of Crypto ’86, volume 263 of Lecture Notes in Computer Science, pages 239–247. Springer, 1987.

[BC86] G. Brassard and C. Crépeau. Non-transitive transfer of confidence: A perfect zeroknowledge interactive protocol for SAT and beyond. In 27^{th} Symp. of Found. of Computer Sci., pages 188–195. IEEE, 1986.

[BCR86] G. Brassard, C. Crépeau, and J.–M. Robert. Information theoretic reductions among disclosure problems. In 27^{th} Symp. of Found. of Computer Sci., pages 168–173. IEEE, 1986.

[Cr86b] C. Crépeau. A secure poker protocol that minimizes the effects of player coalitions. In Advances in Cryptology: Proceedings of Crypto ’85, vol 218 of LNCS,pages 73–86. Springer, 1986.